Two coins tossed simultaneously.A person receive Rs.8 each head and losses Rs.10 for each tail.Find expected value of the amount


1
Expert's answer
2021-01-18T15:10:08-0500

Two coins tossed simultaneously. Let "S" be the sample space .

Then "S""=" {"HH,HT,TH,TT" }

Let us define a random variable "X:S\to R" is defined by

"X(a)=" Net amount gain , "\forall a\in S"

Then, "X(HH)=16,X(HT)=-2,X(TH)=-2,X(TT)=-20" .

And probability of each point "X" is given below,

"P(X=16)=\frac{1}{4},P(X=-2)=\frac{1}{2},P(X=-20)=\frac{1}{4}"

Now let us draw a table of distribution function as follows




Then "E(X)=\Sigma x_i.P(x_i)"

"=(16×\frac{1}{4})+(-2×\frac{1}{2})+(-20×\frac{1}{4})"

"=(4-1-5)" "=-2"

Therefore expected value of amount is "(-2)."